18% silver alloy is made using a combination of 14% silver alloy and 24% silver alloy. How much of each alloy is used to make 20 oz of 18% silver alloy?
Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

.18*20=.14x+.24*(20-x)
x=12 oz at 14%
20-x=8 oz at 24%

Answer by Alan3354(69443)

18% silver alloy is made using a combination of 14% silver alloy and 24% silver alloy. How much of each alloy is used to make 20 oz of 18% silver alloy?
-----------
0.14x + 0.24(20-x) = 0.18*20
14x + 480-24x = 360
-10x = -120
x = 12 oz of 14%
---> 8 oz of 20%